### v1

@@ Abstract
This paper introduces a method that globally converges to B-stationary points of mathematical programs with equilibrium constraints (MPECs) in a finite number of iterations. B-stationarity is necessary for optimality and means that no feasible first-order direction improves the objective. Given a feasible point of an MPEC, B-stationarity can be certified by solving a linear program with equilibrium constraints (LPEC) constructed at this point. The proposed method solves a sequence of LPECs, which either certify B-stationarity or provide an active-set estimate for the complementarity constraints, and nonlinear programs (NLPs) – referred to as branch NLPs (BNLPs) – obtained by fixing the active set in the MPEC. A BNLP is more regular than the MPEC, easier to solve, and with the correct active set, its solution coincides with the solution of the MPEC. The method has two phases: the first phase identifies a feasible BNLP or certifies local infeasibility, and the second phase solves a finite sequence of BNLPs until a B-stationary point of the MPECis found. The paper provides a detailed convergence analysis and discusses implementation details. In addition, extensive numerical experiments and an open-source software implementation are provided. The experiments demonstrate that the proposed method is more robust and faster than relaxation-based methods, while also providing a certificate of B-stationarity without requiring the usual too-restrictive assumptions.
Keywords:
MPEC MPCC nonlinear programming complementarity constraints equilibrium constraints
MSC:
90C30 90C33 49M37 65K1090C11
@@ 1 Introduction
with the partition of variables $x=(x_{0},x_{1},x_{2})\in{\mathbb{R}}^{n}$, with $x_{0}\in{\mathbb{R}}^{p}$, $x_{1},x_{2}\in{\mathbb{R}}^{m}$. The functions $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$ and $c:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n_{c}}$ are assumed to be at least twice continuously differentiable. The notation $0\leq x_{1}\perp x_{2}\geq 0$, means that all components for the vectors $x_{1}$ and $x_{2}$ must be nonnegative, and that vectors $x_{1}$ and $x_{2}$ are orthogonal, i.e., either $x_{1,i}=0$ or $x_{2,i}=0$ for all $i$. To keep the notation lighter we do not explicitly include the equality constraint in (1). However, if they are defined via at least twice continuously differentiable functions, all results in this paper are readily extended to cases that include equality constraints.
Strictly speaking, equilibrium constraints are more general than complementarity constraints (1c), and they are equivalent under suitable conditions Luo1996 . In this paper, we focus on mathematical programs with complementarity constraints, but as common in the optimization literature, we stick to the easier-to-pronounce acronym MPEC.
Solving MPECs reliably and fast is of great practical interest, as they arise in numerous applications including process engineering Baumrucker2008 , robotics Wensing2023 , optimal control of hybrid dynamical systems Nurkanovic2023f ; Nurkanovic2022b , bi-level optimization Kim2020 , nonsmooth optimization Hegerhorst2020 , inverse optimization Albrecht2017 ; Hu2012 , to name a few examples. A related class of problems of interest, equivalent to MPECs Achtziger2008 , are mathematical programs with vanishing constraints.
@@
However, this NLP is degenerate from a numerical and theoretical point of view. In particular, standard constraint qualifications such as the Mangasarian-Fromovitz constraint qualification (MFCQ) (and thus the stronger linear independence constraint qualification (LICQ)) are violated at all feasible points of this NLP Jane2005 ; Scheel2000 . Consequently, standard NLP methods applied to (2) can become quite inefficient Kim2020 ; Nurkanovic2024b , because the set of its Lagrange multipliers is unbounded Gauvin1977 . Thus, the standard Karush-Kuhn-Tucker (KKT) conditions may not be directly applicable to (2), which complicates the definition and certification of a stationary point Scheel2000 . There exist many other reformulations of the complementarity constraints (1c), including smooth and nonsmooth C-functions Facchinei2003 ; Leyffer2006a , but none of them circumvents the inherent degeneracy of NLP reformulations of MPECs, caused by the complementarity constraints.
A necessary condition for optimality is that at the local minimizer, there are no feasible first-order descent directions (formal definitions are given in the next section). Points that satisfy this condition are called stationary points (local optimality is further certified via sufficient conditions). In the MPEC literature, such points are called B-stationary points. For standard regular NLPs, B-stationary points are simply points that satisfy the KKT conditions. However, directly solving the KKT conditions of the NLP (2) might be too difficult or impossible because of its degeneracy. The MPEC literature has introduced many weaker stationarity notions, of which most may allow first-order descent directions, making them a too weak stopping criterion for a practical algorithm Kirches2022 ; Leyffer2007 . This is further discussed in Section 2.
@@
Regularization methods replace the difficult constraints  (2c) with a smoothed or relaxed version of them. For example, in Scholtes’ global relaxation Scholtes2001 $x_{1,i}x_{2,i}\leq 0$ is replaced by $x_{1,i}x_{2,i}\leq\tau$ for all $i$ and $\tau>0$. This results in a regular NLP that usually satisfies MFCQ and can be solved by existing NLP algorithms. In practice, a sequence of such problems with a decreasing $\tau$ is solved to get a good approximation of the MPEC solution. There are many sophisticated variations for replacing (2c) by a more regular set of constraints DeMiguel2005 ; Kadrani2009 ; Kanzow2013 ; Lin2003 ; Scholtes2001 ; Steffensen2010 . An alternative are penalty methods Anitescu2005 ; Anitescu2005a ; Ralph2004 , that treat the complicating constraints $x_{1,i}x_{2,i}\leq 0$ by adding a penalty to the objective function of (2), e.g. $\rho\sum_{i=1}^{m}x_{i,1}x_{2,i}$. Then a sequence of problems with increasing $\rho>0$ is solved to compute a solution approximation. Under appropriate conditions, even for finite $\rho$, a stationary point of an MPEC can be computed by solving a standard NLP Anitescu2005 ; Ralph2004 . Interestingly, under appropriate assumptions, there is a one-to-one correspondence between the iterates of some relaxation and penalty methods Leyffer2006 . However, unless very restrictive assumptions hold, all these methods may converge to stationary points weaker than B-stationarity Hoheisel2013 ; Kanzow2015 .
@@
Active set methods take a different approach to obtaining a sequence of regular NLPs, whose solutions eventually solve the MPEC. The main idea is to fix the active set complementarity constraints, i.e., replace (1c) by $x_{1,i}=0,x_{2,i}\geq 0,\forall i\in{\mathcal{I}}_{1}$ and $x_{1,j}\geq 0,x_{2,j}\geq 0,\forall j\in{\mathcal{I}}_{2}$, where ${\mathcal{I}}_{1}$ and ${\mathcal{I}}_{2}$ are a partition of $\{1,\ldots,m\}$. The resulting NLPs – called branch NLPs (BNLPs) – are now regular problems that can be solved (approximately) with classical NLP methods. Given the correct partition, the solution of such an NLP also solves the MPEC (1). However, finding such a partition and certifying optimality has combinatorial complexity Scheel2000 . Most active set methods assume that MPEC-LICQ holds (cf. Def. 2.2) and search for appropriate partitions by looking at the Lagrange multipliers of the active complementarity constraints to either find a new active set or to certify B-stationarity. Such methods are reported in Fukushima2002 ; Giallombardo2008 ; Izmailov2008 ; Jiang1999 ; Lin2006 ; Liu2006 ; Luo1996 ; Scholtes1999 . However, if MPEC-LICQ does not hold, only weaker concepts of stationarity – possibly allowing descent directions – can be used as stopping criteria.
@@
To avoid the restrictive assumption of MPEC-LICQ, B-stationarity can be verified by solving a linear program with equilibrium constraints (LPEC) Luo1996 , obtained by linearizing the functions $f$ and $c$ in (1), but retaining the complementarity constraints. If the current point is not B-stationary then the LPEC provides a descent direction and a corresponding active set guess. Such an approach is taken in Guo2022 ; Kirches2022 ; Leyffer2007 . Leyffer and Munson Leyffer2007 propose a sequential LPEC quadratic programming (QP) method. Since solving an LPEC can be computationally expensive, Kirches et al. Kirches2022 consider easy-to-solve LPECs that have only complementarity and box constraints. They suggest treating the remaining equality and inequality constraints by an augmented Lagrangian formulation. Guo and Deng study such an approach Guo2022 , where convergence to M-stationary points (cf. Def. 2.3) is proved. However, M-stationarity may allow first-order descent directions Leyffer2007 . The main practical drawback of these methods is the difficulty of implementation, since all the ingredients necessary for robustness must be provided, e.g. promotion of global convergence, treatment of indefinite Hessians, second-order corrections, and efficient subproblem solvers.
@@
Finally, we mention the class of hybrid methods Kazi2024 ; Lin2006 , which combine active set and regularization-based methods. The method presented here, which we call MPEC optimizer (MPECopt), can be classified as such a method. It consists of two phases. The goal of Phase I is to find a first feasible BNLP. We propose two distinct realizations for Phase I. First, we propose to solve a sequence of relaxed MPECs to find solution approximations, and LPECs to identify a valid active set corresponding to a feasible BNLP. Second, to find a feasible BNLP, we propose to solve an easier auxiliary MPEC, where the objective is to minimize the constraint violation, and the only constraints are the complementarity constraints. This problem is always feasible and can be directly treated with Phase II. The second phase starts from a feasible BNLP and solves LPECs constructed at the local minimizer of the current BNLP. The LPEC solution either certifies B-stationarity or provides an active set guess for a BNLP whose solution has a better objective value. Notably, the LPECs steps are never applied to update the current iterate, but only to provide a better active set guess or to certify B-stationarity.
@@
We comment on the differences to other hybrid approaches in the literature. In contrast to Lin and Fukushima’s method Lin2006 , our proposed method does not use multipliers but solves LPECs for a new active set guess, and thus does not require MPEC-LICQ. As a consequence, our method, unlike the method in Lin2006 , does not preclude convergence to B-stationary points that are not S-stationary. In contrast to Kazi et al. Kazi2024 , we never apply the LPEC steps. This leads to several advantages: all iterates of phase II are feasible, finite termination is ensured, and the BNLPs are often easier to solve than relaxed MPEC subproblems.
In summary, having an NLP and LPEC solver at hand, the proposed method is easy to implement and finds B-stationary points in a finite number of steps without the usual restrictive assumptions. LPECs have due to the complementarity constraints also a combinatorial difficulty. Fortunately, as we demonstrated in numerous experiments, LPECs can be solved efficiently in practice either with regularization-based or mixed-integer linear programming (MILP) methods Gurobi ; Huangfu2018 . Furthermore, we discuss the theoretical reasons why this is the case. We mention that also dedicated LPEC methods exist Hu2012 ; Fang2010 ; JaraMoroni2018 ; JaraMoroni2020 , which could be used within the proposed algorithm. All algorithmic details are discussed in Section 3.
Notation: In this paper, subscripts are used to identify components of vectors or matrices, e.g. $x_{i}$, and superscripts to denote iterates, e.g. $x^{k}$. Functions that are evaluated at particular iterates are denoted as $f^{k}:=f(x^{k})$. The concatenation of two vectors $a\in R^{n}$ and $b\in{\mathbb{R}}^{m}$ is shortly written as $(a,b)=\begin{bmatrix}a^{\top}&b^{\top}\end{bmatrix}^{\top}$. The same notation is adapted accordingly to the concatenation of several vectors. Given two scalars $a,b\in{\mathbb{R}}$, $\max(a,b)$ returns the larger of them. Let $a,b\in{\mathbb{R}}^{n}$, then $\max(a)$ returns the largest component of $a$, and $c=\max(a,b)$ is a vector in ${\mathbb{R}}^{n}$ with the components $c_{i}=\max(a_{i},b_{i}),i=1,\ldots,n$. The $\min$ function is defined in all cases accordingly. Let $x\in{\mathbb{R}}^{n}$, then $\mathrm{diag}(x)\in{\mathbb{R}}^{n\times n}$ returns a diagonal matrix with $x_{i}$ as its diagonal elements.
@@
Contributions: This paper presents a new method, called MPEC optimizer (MPECopt) that computes B-stationary points by solving a finite number of LPECs and NLPs. The LPECs correspond to the definition of B-stationarity. In our method, they additionally have a trust-region constraint on the step lengths. We discuss in detail two possible different LPEC formulations and their relation, and strengths and weaknesses. Remarkably, at feasible points of the MPEC, the trust-region radius in the LPEC can be arbitrarily small, leading to tight relaxations of the MILP equivalent to the LPEC. This leads to efficient solutions in practice. The proposed algorithm consists of two phases. The first phase finds a feasible BNLP or certifies local infeasibility. We discuss in detail two different algorithms for Phase I, one is based on regularization methods, and the other solves a feasibility problem. We prove that both are guaranteed to find a feasible point or to certify local infeasibility. Phase II iterates over the feasible BNLPs while strictly decreasing the objective until a B-stationary point is found. We prove that Phase II finds a B-stationary point in a finite number of steps. Furthermore, we compare the new method with several relaxation-based methods on the MacMPEC and a synthetic nonlinear MPEC benchmark. Higher success rates and generally faster computation times are reported. An open-source prototype implementation of the method is provided at https://github.com/nosnoc/mpecopt.
@@
Outline: Section 2 reviews the first-order optimality conditions for MPECs and some standard solution methods. Section 3 provides a detailed statement of MPECopt, discusses how to formulate and solve LPECs, and provides all implementation details. Section 4 gives a detailed convergence analysis of the new method. Section 5 provides extensive numerical results on several benchmark test sets. Finally, Section 6 concludes the paper and discusses future research directions.

### v2

@@ Abstract
This paper introduces a computationally efficient method that globally converges to B-stationary points of mathematical programs with equilibrium constraints (MPECs) in a finite number of iterations. B-stationarity is necessary for optimality and means that no feasible first-order direction can improve the objective. Given a feasible point of an MPEC, B-stationarity can be certified by solving a linear program with equilibrium constraints (LPEC) constructed at this point. The proposed method solves a sequence of LPECs, which either certify B-stationarity or provide an active-set estimate for the complementarity constraints, along with nonlinear programs (NLPs) – referred to as branch NLPs (BNLPs) – obtained by fixing the active set in the MPEC. A BNLP is more regular than the original MPEC, easier to solve, and with the correct active set, its solution coincides with that of the MPEC. We show that, unless the current iterate is B-stationary, these combinatorial LPECs need not be solved to optimality; for convergence, it suffices to compute a nonzero feasible point, yielding significant computational savings. The method proceeds in two phases: the first identifies a feasible BNLP or a stationary point of a constraint infeasibility minimization problem, and the second solves a finite sequence of BNLPs until a B-stationary point of the MPECis found. We established finite convergence under the MPEC-MFCQ. Numerical experiments and an open-source software implementation show that the proposed method is more robust and faster than relaxation-based and mixed-integer NLP approaches, even on medium to large-scale instances.
@@ 1 Introduction
with the partition of variables $x=(x_{0},x_{1},x_{2})\in{\mathbb{R}}^{n}$, with $x_{0}\in{\mathbb{R}}^{p}$, $x_{1},x_{2}\in{\mathbb{R}}^{m}$. The functions $f:{\mathbb{R}}^{n}\to{\mathbb{R}}$ and $c:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n_{c}}$ are assumed to be at least twice continuously differentiable. The notation $0\leq x_{1}\perp x_{2}\geq 0$ means that $x_{1,i},x_{2,i}\geq 0$ and $x_{1,i}x_{2,i}=0$ for all $i$. There is no loss of generality in using (1c), since the more general form $0\leq g(x)\perp h(x)\geq 0$, with $g,h:{\mathbb{R}}^{p}\to{\mathbb{R}}^{m}$ twice continuously differentiable, can be transformed into (1c) by introducing slack variables $x_{1}=g(x)$ and $x_{2}=h(x)$. For lighter notation we omit equality constraints in (1). If given by twice continuously differentiable functions, all results extend directly to this case.
Equilibrium constraints are more general than complementarity constraints (1c), but coincide under suitable conditions Luo1996 . We focus on complementarity constraints, yet follow the literature in using the acronym MPEC.
Solving MPECs reliably and fast is of great practical interest, as they arise in numerous applications including process engineering Baumrucker2008 , robotics Wensing2023 , optimal control of hybrid dynamical systems Nurkanovic2023f ; Nurkanovic2022b , bi-level optimization Kim2020 , nonsmooth optimization Hegerhorst2020 , inverse optimization Albrecht2017 ; Hu2012 , to name a few examples. A related class of problems of interest, equivalent to MPEC, are mathematical programs with vanishing constraints Achtziger2008 .
@@
However, this NLP is numerically and theoretically degenerate. Standard constraint qualifications such as the Mangasarian-Fromovitz constraint qualification (MFCQ) fail at all feasible points Jane2005 ; Scheel2000 , making standard NLP methods inefficient Kim2020 ; Nurkanovic2024b due to unbounded Lagrange multiplier sets Gauvin1977 . Moreover, standard KKT conditions may not apply to (2), complicating stationary point definition and certification Scheel2000 . Various reformulations of (1c) exist, including smooth and nonsmooth C-functions Facchinei2003 ; Leyffer2006a , but none resolve the inherent degeneracy of NLP reformulations of MPECs.
A necessary optimality condition is the absence of feasible first-order descent directions. In the MPEC literature, such points are called B-stationary points. For regular NLPs, B-stationary points simply satisfy KKT conditions, but directly solving KKT conditions for (2) may be difficult or impossible due to degeneracy. Weaker stationarity notions exist, but most allow first-order descent directions, making them too weak as practical stopping criteria Leyffer2007 .
@@
Regularization methods replace the difficult constraints  (2c) with a smoothed or relaxed version of them. For example, in Scholtes’ global relaxation Scholtes2001 $x_{1,i}x_{2,i}\leq 0$ is replaced by $x_{1,i}x_{2,i}\leq\tau$ for all $i$ and $\tau>0$. This yields a regular NLP that usually satisfies MFCQ and can be solved by standard NLP algorithms. In practice, a sequence of NLPs with decreasing $\tau$ is solved to approximate the MPEC solution. Many refinements of this idea exist DeMiguel2005 ; Kadrani2009 ; Kanzow2013 ; Lin2003 ; Scholtes2001 ; Steffensen2010 . Penalty methods Anitescu2005 ; Anitescu2005a ; Ralph2004 handle $x_{1,i}x_{2,i}\leq 0$ by adding a penalty term to the objective, e.g. $\rho\sum_{i=1}^{m}x_{1,i}x_{2,i}$. A sequence of NLPs with increasing $\rho>0$ is then solved to approximate a solution. Under suitable conditions, even finite $\rho$ may yield an MPEC stationary point via a standard NLP Anitescu2005 ; Ralph2004 . However, unless restrictive assumptions hold, all these methods may converge to points that are not B-stationarity Hoheisel2013 ; Kanzow2015 .
@@
Active set methods take a different approach to obtaining a sequence of regular NLPs, whose solutions eventually solve the MPEC. The main idea is to fix the active set complementarity constraints, i.e., replace (1c) by $x_{1,i}=0,x_{2,i}\geq 0,\forall i\in{\mathcal{I}}_{1}$ and $x_{1,j}\geq 0,x_{2,j}=0,\forall j\in{\mathcal{I}}_{2}$, where ${\mathcal{I}}_{1}$ and ${\mathcal{I}}_{2}$ are a partition of $\{1,\ldots,m\}$. The resulting NLPs – called branch NLPs (BNLPs) – are now regular problems that can be solved with classical NLP methods. Given the correct partition, the solution of such an NLP also solves the MPEC (1). However, finding such a partition and certifying optimality has combinatorial complexity Scheel2000 . Most active set methods assume that MPEC-LICQ holds (cf. Def. 2.2) and search for partitions by looking at the Lagrange multipliers of the active complementarity constraints to either find a new active set or to certify B-stationarity. Such methods are reported in Fukushima2002 ; Giallombardo2008 ; Izmailov2008 ; Jiang1999 ; Lin2006 ; Liu2006 ; Luo1996 ; Scholtes1999 . However, if MPEC-LICQ does not hold, only weaker concepts of stationarity – possibly allowing descent directions – can be used as stopping criteria.
@@
To avoid the restrictive assumption of MPEC-LICQ, B-stationarity can be verified by solving a linear program with equilibrium constraints (LPEC) Luo1996 , obtained by linearizing the functions $f$ and $c$ in (1), but retaining the complementarity constraints. If the current point is not B-stationary then the LPEC provides a descent direction and a corresponding active set guess. Such an approach is taken in Guo2022 ; Kirches2022 ; Leyffer2007 . Leyffer and Munson Leyffer2007 propose a sequential LPEC QP method. Since solving an LPEC can be expensive, Kirches et al. Kirches2022 study simpler LPECs with only complementarity and box constraints, and suggest handling of other constraints via an augmented Lagrangian. Guo and Deng Guo2022 analyze this approach and prove convergence to M-stationary points (Def. 2.3). However, M-stationarity may still admit descent directions, cf. (Leyffer2007, , Sec 2.1). A main drawback of these methods is their implementation complexity, since all components of a robust solver must be provided.
@@
Finally, we mention the class of hybrid methods Kazi2024 ; Lin2006 , which combine active set and regularization-based methods. The method presented here, which we call MPEC optimizer (MPECopt), can be classified as such a method and consists of two phases. Phase I finds a feasible BNLP by solving relaxed MPECs for solution approximations, then LPECs to identify valid active sets. Phase II starts from a feasible BNLP and solves LPECs at each BNLP’s stationary point until reaching B-stationarity. LPEC solutions either certify B-stationarity or provide active set guesses for BNLPs with better objective values. Notably, the LPECs steps never update iterates, but only provide better active sets or certify B-stationarity.
@@
We highlight key differences from other hybrid approaches. Unlike Lin and Fukushima Lin2006 , our method solves LPECs for active set guesses rather than using multipliers, eliminating MPEC-LICQ requirements and allowing convergence to B-stationary points that are not S-stationary. Unlike Kazi et al. Kazi2024 , we never apply LPEC steps directly, ensuring all Phase II iterates remain feasible and guaranteeing finite termination.
It is important to note that LPECs are nonconvex combinatorial problems whose efficient solution is crucial for our method’s effectiveness. LPECs with only bound and complementarity constraints have linear complexity (Kirches2022, , Propositions 2.3 and 2.4). As special cases of MPECs, LPECs can be solved with regularization and penalty methods. Tailored LPEC methods include simplex extensions Fang2010 , mixed-integer linear programming (MILP) approaches with tailored cuts Hu2012 ; JaraMoroni2020 , and specialized penalty methods Hall2024 ; JaraMoroni2018 ; see Judice2012 for a survey. However, regularization and penalty methods may converge to points with trivial descent directions and fail to verify B-stationarity, making combinatorial methods preferable. Importantly, we show that, in our proposed method, LPECs need only be solved to optimality when certifying a B-stationary point at the last iteration of Phase II. Otherwise, in Phase I, it is sufficient to find a feasible point of the LPEC, and in Phase II, a feasible descent direction. Our numerical experiments show that this is often equivalent to solving a single LP.
Contributions: This paper presents a new method, called MPEC optimizer (MPECopt) that computes B-stationary points by solving a finite number of LPECs and NLPs. In our method, the LPECs additionally have a trust-region constraint on the step lengths. We discuss in detail two possible different LPEC formulations and their relation.
We show that, unless we need to certify B-stationarity, it is sufficient to find a feasible (improving) point of the LPEC to ensure progress. Moreover, we discuss special cases when LPECs are easy to solve, and cases where a MILPs reformulation can be solved efficiently.
@@
We prove that, by finding a feasible point of an LPEC constructed at a point sufficiently close to the feasible set of the MPEC, the active set obtained from the LPEC corresponds to a BNLP that is guaranteed to be feasible. In practice, this helps to find feasible BNLPs in Phase I by combining well-established regularization methods and an LPEC solver. Phase II iterates over the feasible BNLPs while strictly decreasing the objective until a B-stationary point is found. We prove that, under the MPEC-MFCQ, Phase II finds a B-stationary point in a finite number of steps. Furthermore, we compare the new method with several relaxation-based methods, and a Mixed-Integer Nonlinear Programming (MINLP) reformulation, on the MacMPEC and a synthetic large-scale nonlinear MPEC benchmark. Higher success rates and generally faster computation times are reported, and solving LPECs turns out not to be a computational bottleneck, even for large problem instances. An open-source implementation of the method is provided at github.com/nosnoc/mpecopt.
Notation: Subscripts denote vector/matrix components ($x_{i}$), superscripts denote iterates ($x^{k}$). Functions evaluated at particular iterates are denoted as $f^{k}:=f(x^{k})$. The concatenation of two vectors $a\in R^{n}$ and $b\in{\mathbb{R}}^{m}$ is shortly written as $(a,b)=\begin{bmatrix}a^{\top}&b^{\top}\end{bmatrix}^{\top}$. The same notation is adapted accordingly to the concatenation of several vectors. For scalars $a,b\in{\mathbb{R}}$, $\max(a,b)$ returns the larger value. For vectors $a,b\in{\mathbb{R}}^{n}$: $\max(a)$ returns the largest component, $c=\max(a,b)$ gives $c_{i}=\max(a_{i},b_{i})$. The $\min$ function is defined in all cases accordingly. Let $x\in{\mathbb{R}}^{n}$, then $\mathrm{diag}(x)\in{\mathbb{R}}^{n\times n}$ returns a diagonal matrix with $x_{i}$ as its diagonal elements.
@@
Outline: Section 2 recalls some preliminaries. Section 3 provides a detailed statement of MPECopt, discusses how to formulate and solve LPECs. Section 4 gives a detailed convergence analysis of the new method. Section 5 provides all implementation details and numerical results on two benchmark test sets. Finally, Section 6 provides concluding remarks.